Let G=D_8, and let N={e,a^2,a^4,a^6}.
(a) Find all left cosets and all right cosets of N, and verify that N is a normal subgroup of G.
(b) Show that G/N has order 4, but is not cyclic.
1
Expert's answer
2020-08-10T15:38:22-0400

"G=D_8=\{e, a, a^2, a^3, x, ax\equiv xa^3, a^2x, a^3x\equiv xa\}\\
a\text{ --- rotation by angle of }\pi/2\text{ counterclockwise}\\
x\text{ --- reflection about the diagonal joining vertices "2" and "4"}\\
ax\text{ --- reflection about the line joining midpoints of }\\
\text{opposite sides "14" and "23"}\\
a^2x\text{ --- reflection about the diagonal joining vertices "1" and "3"}\\
a^3x\text{ --- reflection about the line joining midpoints of }\\
\text{opposite sides "12" and "34"}\\
N=\{e, a^2, a^4\equiv e, a^6\equiv a^2\}=\{e, a^2\}\text{ --- }\\
\text{subgroup of G}\\
(a)\text{ Left cosets:}\\
1) e\in G\\
eN\equiv N\\
2) a\in G\\
aN=\{a, a^3\}\\
3) a^2\in G\\
a^2N=\{a^2, a^4\}\\
4) a^3\in G\\
a^3N=\{a^3, a^5\equiv a\}\\
5) x\in G\\
xN=\{x, xa^2\equiv a^2x\}\\
6) ax\in G\\
(ax)N=\{ax, (ax)a^2\equiv (xa^3)a^2\equiv x(a^3a^2)\equiv xa\}\\
7) a^2x\in G\\
(a^2x)N=\{a^2x, (a^2x)a^2\equiv (xa^2)a^2\equiv x(a^2a^2)\equiv x\}\\
8) xa\in G\\
(xa)N=\{xa, (xa)a^2\equiv xa^3\equiv ax\}\\
\text{Right cosets:}\\
1) e\in G\\
Ne\equiv N\\
2) a\in G\\
Na=\{a, a^3\}\\
3) a^2\in G\\
Na^2=\{a^2, a^4\}\\
4) a^3\in G\\
Na^3=\{a^3, a^5\equiv a\}\\
5) x\in G\\
Nx=\{x, a^2x\}\\
6) ax\in G\\
N(ax)=\{ax, a^2(ax)\equiv (a^2a)x\equiv xa\}\\
7) a^2x\in G\\
N(a^2x)=\{a^2x, a^2(a^2x)\equiv x\}\\
8) xa\in G\\
N(xa)=\{xa, a^2(xa)\equiv a^2(a^3x)\equiv ax\}\\
\text{We can see that the corresponding left and right}\\
\text{cosets are equal. So using the criterion of a normal}\\
\text{subgroup we can conclude that }N\text{ is a normal}\\
\text{subgroup of G}.\\
(b) G/N \text{ has order } \frac{|G|}{|N|}=\frac{8}{2}=4 \text{ (using Lagrange's theorem)}\\
 G/N \text{ is a group of all the left cosets }\\
 G/N \text{ is not cyclic because there is no element } a \text{ such that } gN=a^i"